# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v2

@@ H.2 Evaluation Metrics
GPT3.5 (4-shot) $61.25$ $48.95$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 39.20}$) $62.72$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 45.22}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.69}$) $28.63$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 12.50}$) $60.07$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 5.80}$)
GPT3.5 (CoT) $66.55$ $66.62$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 21.53}$) $96.85$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 79.35}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.69}$) $50.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 9.68}$) $78.57$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.30}$)
GPT3.5 (SG) (Our) $70.61$ $73.30$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 14.86}$) $92.93$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 75.43}$) $99.22$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 28.91}$) $49.19$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 8.06}$) $78.66$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.39}$)
@@
GPT4o (4-shot) $66.23$ $67.87$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.40}$) $98.70$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.11}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00}$) $78.63$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 8.06}$) $86.30$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.14}$)
GPT4o (CoT) $72.44$ $72.77$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.51}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.20}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00}$) $73.79$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.23}$) $86.64$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.48}$)
GPT4o (SG) (Our) $76.44$ $74.67$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.40}$) $97.72$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 1.09}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00}$) $68.55$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 2.02}$) $85.23$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.08}$)
@@
Figure 12: Red shows the wrong FoR identifications, and green shows the correct ones. The dark color is for relative FoRs, while the light color is for intrinsic FoRs. The round shape is for the external FoRs, while the square is for internal FoRs. The depth of the plots shows the four FoRs, i.e., external relative, external intrinsic, internal intrinsic, and internal relative, from front to back.
@@
Figure 13: Red shows the wrong FoR identifications, and green shows the correct ones. The dark color is for relative FoRs, while the light color is for intrinsic FoRs. The round shape is for the external FoRs, while the square is for internal FoRs. The depth of the plots shows the four FoRs, i.e., external relative, external intrinsic, internal intrinsic, and internal relative, from front to back. This plot is the result of the rest of LLMs.
H.3 Results
H.3.1 FoR Inherently Bias in LLMs
@@
C-spilt. The zero-shot setting reflects the LLMs’ inherent bias in identifying FoR. Table 13 presents the accuracy for each FoR class in C-split, where sentences explicitly include information about topology and perspectives. We found that some models strongly prefer specific FoR classes. Notably, Gemme2-9B achieves a near-perfect accuracy on external relative FoR but performs poorly on other classes, especially external intrinsic, indicating a notable bias towards external relative. In contrast, GPT4o and Qwen2-72B perform exceptionally in both intrinsic FoR classes. However, they perform poorly in the relative FoRs. We also
@@
A-spilt. We examine the FoR bias in the A-split. Based on the results in Table 13, we plotted the top-3 models’ results (Gemma2-9B, Llama3-70B, and GPT4o) for a more precise analysis in Figures 12. The plots show the frequencies of each FoR category. According to the plot, Gemma and GPT have strong biases toward external relative and external intrinsic, respectively. This bias helps Gemma2 perform well in the A-split since all spatial expressions can be interpreted as external relative. However, GPT4o’s bias leads to errors when intrinsic FoRs aren’t valid, as in the Box and Pen cases (see plots (c) and (d)). Llama3 exhibits different behavior, showing a bias based on the relatum’s properties, specifically the relatum’s affordance as a container. In cases where relatum cannot serve as containers, i.e., Cow and Pen cases, Llama3 favors external relative. Conversely, Llama3 tends to favor external intrinsic when the relatum has the potential to be a container.
H.3.2 Behavior with ICL variations
@@
C-spilt. We evaluate the models’ behavior under various in-context learning (ICL) methods. As observed in Table 13, the few-shot method improves the performance of the zero-shot method across multiple LLMs by reducing their original bias toward specific classes. Reducing the bias, however, lowers the performance in some cases, such as the performance of Gemma 2 in ER class. One noteworthy observation is that while the CoT prompting generally improves performance in larger LLMs, it is counterproductive in smaller models for some FoR classes. This suggests that the smaller models have difficulty inferring FoR from the longer context. This negative effect also appears in SG prompting, which uses longer explanations. Despite performance degradation in particular classes of small models, SG prompting performs exceptionally well across various models and achieves outstanding performance with Qwen2-72B. We further investigate the performance of CoT and SG prompting. As shown in Table 12, CoT exhibits a substantial difference in performance between contexts with inherently clear FoR and contexts requiring the template to clarify FoR ambiguity. This implies that CoT heavily relies on the specific template to identify FoR classes. In contrast, SG prompting demonstrates a smaller gap between these two scenarios and significantly enhances performance over CoT in inherently clear FoR contexts. Therefore, guiding the model to provide characteristics regarding topological, distance, and directional types of relations improves FoR comprehension.
@@
A-spilt. We use the same Figure 12 to observe the behavior when applying ICL. The A-split shows minimal improvement with ICL variations, though some notable changes are observed. With few-shot, all models show a strong bias toward external intrinsic FoR, even when the relatum lacks intrinsic directions, i.e., Box and Pen cases. This bias appears even in Gemma2-9B, which usually behaves differently. This suggests that the models pick up biases from the examples despite efforts to avoid such patterns. However, CoT reduces some bias, leading LLMs to revisit relative, which is generally valid across scenarios. In Gemma2, the model predicts relative FoR where the relatum has intrinsic directions, i.e., Cow and Car cases. Llama3 behaves similarly in cases where the relatum cannot act as a container, i.e., Cow and Pen cases. GPT4o, however, does not depend on the relatum’s properties and shows slight improvements across all cases. Unlike CoT, our SG prompting is effective in all scenarios. It significantly reduces biases while following a similar pattern to CoT. Specifically, SG prompting increases external relative predictions for Car and Cow in Gemma2-9B, and for Cow and Pen in Llama3-70B. Nevertheless, GPT4o shows only a slight bias reduction. However, Our proposed method improves the overall performance of most models, as shown in Table 13. The Llama3-70B behaviors are also seen in LLama3-8B and GPT3.5. The plots for LLama3-8B and GPT3.5 are in Figure 13.
H.3.3 Experiment with different temperatures
@@
We conducted additional experiments to further investigate the impact of temperature on the biased interpretation of the model in the A-split of our dataset. As presented in Table 14, comparing distinct temperatures (0 and 1) revealed a shift in the distribution. The frequencies of the classes experienced a change of up to 10%. However, the magnitude of this change is relatively minor, and the relative preferences for most categories remained unchanged. Specifically, the model exhibited the highest frequency responses for the cow, car, and pen cases, even with higher frequencies in certain settings. Consequently, a high temperature does not substantially alter the diversity of LLMs’ responses to this task, which is an intriguing finding.
@@
Table 14: The results between two different temperatures of Llam3-70B on the A-spilt of FoREST. The number shows the percentage frequency of responses from the model.
Appendix I Study of Human bias in A-spilt
I.1 Experiment Setting
We provide a Google Form with instructions in Listing 1, followed by 150 randomly ordered context-question pairs. These are sampled based on relatum category: 25 questions per case from the A-split, covering four cases (Cow, Car, Box, Pen) for camera-perspective questions, and two cases (Cow, Car) for relatum-perspective questions, as Box and Pen lack intrinsic direction. No time constraints were imposed, and all participants were compensated at the standard research assistant rate. All participants were informed about the study details at the end and asked for consent to use their responses in the analysis. After acquiring the results, we utilize the same metric defined in Section 5.1. Then, we average the accuracy of each participant to derive the human baseline based on this study.
I.2 Results
Individual human results are shown in Table 15. These results indicate that humans rely heavily on their assumptions when interpreting ambiguous frame-of-reference (FoR) contexts. Most participants favored the intrinsic FoR in A-split scenarios, though this preference varied across individuals, suggesting that background or prior visual experience may influence FoR interpretation. This aligns with findings from cognitive studies on how humans describe spatial relations that recently seen examples may influence Coventry et al. (2018). However, our results contrast with other studies suggesting that humans tend to adopt an egocentric (relative) FoR in scenes with fewer landmarks and an allocentric (intrinsic) FoR in those with more landmarks Forster et al. (2023). Since our study does not provide visual input, we cannot infer how participants mentally reconstruct scenes to answer the questions with only two objects in the scene. The next notable point is that while Box and Pen cases appear straightforward for models (Table 1), humans may still struggle with questions in these categories. Humans do not exhibit the same limitations as LLMs in perspective-shifting tasks. When the context and question differ in perspective, humans still achieve over 90% accuracy, whereas LLMs often struggle with this type of reasoning.
Instruction:
Camera perspective Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
RIAcc. RIAcc. Acc. Acc. Acc. RIAcc. RIAcc. Acc.
Table 15: QA accuracy in the A-Split of human study. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Appendix J Additional Results with GPT-o4-mini-high
We include results from GPT-4o-mini-high, a model configured for enhanced reasoning, for comparison with our previous results in Table 1 and Table 2. In this experiment, the reasoning-effort parameter is set to high. As shown in Table 16, the model exhibits a balanced preference between relative and intrinsic interpretations. Similar to Qwen2VL, it demonstrates strong reasoning ability, achieving high accuracy even when FoR adaptation is required due to its visual reasoning capability. Notably, GPT-o4-mini-high performs well on questions from the relatum perspective, even outperforms the GPT-4o (SG+CoT) in Table 17. However, its performance declines on camera-perspective questions. We believe this is possibly due to training emphasizing alternative perspectives over the camera view, which may reduce their performance from the camera perspective. Therefore, even with a strong reasoning model, the perspective shift remains challenging in the textual domain.
Camera perspective Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
RIAcc. RIAcc. Acc. Acc. Acc. RIAcc. RIAcc. Acc.
Table 16: Additional results of QA accuracy in the A-Split with GPT-o4-mini-high. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Camera perspective Relatum perspective
GPT-4o (0-shot) $79.7$ $45.1$ $39.5$ $90.2$ $64.2$ $46.9$ $88.5$ $98.2$ $34.8$ $67.5$
Appendix K LLM Prompt
K.1 FoR Identification
@@ I.2 Results
K.2 Question Answering Example

### v3

@@ Appendix I VISOR-uncond Results
GPT3.5 (4-shot) $61.25$ 48.9548.95($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 39.20$}) 62.7262.72($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 45.22$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.69$}) 28.6328.63($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 12.50$}) 60.0760.07($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 5.80$})
GPT3.5 (CoT) $66.55$ 66.6266.62($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 21.53$}) 96.8596.85($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 79.35$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.69$}) 50.8150.81($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 9.68$}) 78.5778.57($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.30$})
GPT3.5 (SG) (Our) $70.61$ 73.3073.30($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 14.86$}) 92.9392.93($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 75.43$}) 99.2299.22($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 28.91$}) 49.1949.19($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 8.06$}) 78.6678.66($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.39$})
@@
GPT4o (4-shot) $66.23$ 67.8767.87($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.40$}) 98.7098.70($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.11$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00$}) 78.6378.63($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 8.06$}) 86.3086.30($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.14$})
GPT4o (CoT) $72.44$ 72.7772.77($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.51$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.20$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00$}) 73.7973.79($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.23$}) 86.6486.64($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 1.48$})
GPT4o (SG) (Our) $76.44$ 74.6774.67($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.40$}) 97.7297.72($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 1.09$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.00$}) 68.5568.55($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 2.02$}) 85.2385.23($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 0.08$})
@@
Figure 12: Red indicates incorrect FoR identifications and green indicates correct ones. Dark colors represent relative FoRs, while light colors represent intrinsic FoRs. Round shapes correspond to external FoRs, and squares correspond to internal FoRs. The plot depth represents the four FoRs—external relative, external intrinsic, internal intrinsic, and internal relative—from front to back. This plot shows the results for Gemma-9B, Llama3-72B, and GPT4o.
@@
Figure 13: Red indicates incorrect FoR identifications and green indicates correct ones. Dark colors represent relative FoRs, while light colors represent intrinsic FoRs. Round shapes correspond to external FoRs, and squares correspond to internal FoRs. The plot depth represents the four FoRs—external relative, external intrinsic, internal intrinsic, and internal relative—from front to back. This plot shows the results for Llama3-8B and GPT3.5.
K.3 Experimental results
K.3.1 Inherent FoR bias in LLMs
@@
C-spilt. The zero-shot setting reflects the LLMs’ inherent bias in identifying FoR. Table 19 shows accuracy for each FoR class in the C-split, where sentences explicitly include topology and perspective information. Some models strongly favor specific FoR classes: notably, Gemma2-9B achieves near-perfect accuracy on external relative FoR but performs poorly on others—especially external intrinsic—indicating a strong bias toward external relative. In contrast, GPT-4o and Qwen2-72B perform well on intrinsic FoR classes but poorly on relative ones.
@@
A-spilt. We examine FoR bias in the A-split. Based on the results in Table 19, we plot the top three models (Gemma2-9B, Llama3-70B, and GPT-4o) for detailed analysis in Figures 12. The plots show the frequency distribution of FoR categories. Gemma2-9B and GPT-4o display strong biases toward external relative and external intrinsic, respectively. This bias benefits Gemma2-9B in the A-split, since all spatial expressions can be interpreted as external relative. By contrast, GPT-4o’s bias leads to errors when intrinsic FoRs are invalid, as in the Box and Pen cases (plots (c) and (d)). Llama3 shows a different pattern, with its bias depending on the relatum’s properties, particularly the container affordance. In cases where the relatum cannot serve as a container (Cow and Pen), Llama3 favors the external relative. Conversely, when the relatum has container potential, Llama3 tends to favor external intrinsic.
K.3.2 Effect of ICL variations
@@
C-spilt. We evaluate model behavior under different in-context learning (ICL) methods. As shown in Table 19, few-shot prompting improves performance over zero-shot across multiple LLMs by reducing their bias toward specific classes, though this reduction sometimes lowers performance (e.g., Gemma2 in the external relative). Another observation is that while CoT generally improves performance in larger LLMs, it can be counterproductive in smaller models for some FoR classes, likely because they struggle to infer FoR from longer contexts. A similar negative effect appears in SG prompting, which also uses longer explanations. Despite these degradations in small models, SG prompting performs well across architectures and achieves outstanding results with Qwen2-72B. To better understand this, we compare CoT and SG prompting in Table 18. CoT shows large performance gaps between contexts with inherently clear FoR and those requiring templates to resolve ambiguity, indicating its reliance on template-specific cues. By contrast, SG prompting exhibits a smaller gap and substantially outperforms CoT in inherently clear contexts. This suggests that guiding models to identify topological, distance, and directional relation characteristics enhances FoR comprehension.
@@
A-spilt. We use Figure 12 to analyze behavior under ICL of the A-split. The A-split shows minimal improvement overall, though some notable changes emerge. With few-shot, all models shift toward external intrinsic FoR—even when the relatum lacks intrinsic direction (Box and Pen)—a bias also observed in Gemma2-9B, which usually behaves differently. This indicates that models inherit biases from examples despite efforts to avoid them. CoT reduces some of this bias, encouraging LLMs to predict relative FoR, which is generally valid across scenarios. For example, Gemma2 predicts relative FoR for Cow and Car, while Llama3 does so for Cow and Pen, where the relatum cannot act as a container. GPT-4o shows slight improvements across all cases without relying on relatum properties. Unlike CoT, SG prompting is effective across scenarios, significantly reducing biases while following a similar adjustment pattern. Specifically, it increases external relative predictions for Car and Cow in Gemma2-9B and for Cow and Pen in Llama3-70B. GPT-4o shows only slight bias reduction, yet overall performance improves for most models (Table 19). Llama3-70B’s behavior is mirrored in Llama3-8B and GPT-3.5, with corresponding plots shown in Figure 13.
K.3.3 Experiment with different temperatures
@@
We conducted additional experiments to investigate the impact of temperature on model bias in the A-split of the FoREST dataset. As shown in Table 20, comparing temperatures 0 and 1 revealed distribution shifts of up to 10%. However, the relative preferences across most categories remain unchanged. In particular, the model produced the highest-frequency responses for the Cow, Car, and Pen cases, with some increases under certain settings. Overall, higher temperature does not substantially increase the diversity of LLM responses in this task, which is a notable finding.
@@
Table 20: Percentage distribution of responses from Llama3-70B at two different temperatures (0 and 1) on the A-split of FoREST. All FoR classes are external relative (ER), external intrinsic (EI), internal intrinsic (II), and internal relative (IR).
Appendix L Prompt Specifications
L.1 FoR identification task
@@
L.2 Question answering task